# Leveraging Priors via Diffusion Bridge for Time Series Generation

### v1

@@ Methods
We now consider trend-guided generation, where the data samples are in $\mathcal{D}=\{({\mathbf{x}}^{i},{\mathbf{y}}^{i})\}_{i=1}^{N}$. For the trend ${\mathbf{y}}$, we consider three types: linear, polynomial with degree of three, and Butterworth filter (Butterworth et al. 1930), a signal processing filter that captures trends (Xia et al. 2024). Using these trends as conditions, we reimplement the experiments in Table 1 with the ETTh and Energy datasets. To apply the condition to baseline methods, we feed the trend into a conditional embedding and concatenate it with the original embedding. For the proposed TimeBridge, we set the prior ${\mathbf{x}}_{T}={\mathbf{y}}$ upon the conditional embedding for translation.
@@
Experimental results are shown in Table 2. The rows with trend baseline indicate the results with the trends without any synthesis. After synthesis with trend conditions, we observe that the quality metrics such as Context-FID, correlation, and discriminative scores diminished compared to the unconditional generation in Table 1, indicating that trend conditions help the model generate more realistic data samples. With the trend condition in ${\mathbf{x}}_{T}$, the proposed TimeBridge achieves the best performance in Context-FID in all settings and the best discriminative score in 5 out of 6 settings while maintaining correlation scores at similar levels to Diffusion-TS. This indicates that the proposed framework can be broadly adopted by different conditioning settings, without controlling any details other than the prior distribution. The sampling path of TimeBridge from trend to data is shown in Figure 4(a).
Imputation with Fixed Masks
@@
For imputation, which is a common setting of hard conditions, we compare the proposed method with imputation diffusion models and the conditional Langevin sampler of Diffusion-TS (Yuan and Qiao 2024). We use random masks for the Mujoco dataset and the geometric mask from Zerveas et al. (2021) for the ETTh and Energy datasets, with a given missing ratio as in Yuan and Qiao (2024). We set the total length of imputation to 100 for the Mujoco dataset, matching the data length, and split the ETTh and Energy datasets into 48 time steps. For the proposed TimeBridge, we preserve the observed values and set the prior ${\mathbf{x}}_{T}$ as the interpolated values of the condition ${\mathbf{c}}$ as in Equation (12).
@@
The imputation results are shown in Table 3 and Table 4. Both tables demonstrate that TimeBridge achieves the lowest measures in imputation tasks. Since the Langevin sampler of Diffusion-TS does not require individual training for various missing ratios, we ensure a fair comparison by evaluating TimeBridge with the same training setting, TimeBridge-75%, which uses 75% missing values for training. The performance of TimeBridge-75% even outperforms the models trained with equal missing values, highlighting the generalization capability of TimeBridge in imputation tasks. The visualization of imputation results is shown in Figure 5, and the sampling path can be found in Figure 4(b).
@@
Figure 5: Visualization of imputation on (Top) ETTh and (Bottom) Energy dataset. The green dotted lines and circles denote true and observed value, respectively. Each method is visualized along with their interval.
@@
Table 5 evaluates the effectiveness of training design with three variants: (1) variance exploding (VE) scheduler, (2) noise matching instead of directly predicting $D_{\theta}$ as in Zhou et al. (2024), and (3) without Fourier-based loss during training. Further experimental results of the ablation study are presented in Appendix D.
@@
Table 5: Ablation study for training and sampling features.
Computation efficiency
Table 6 illustrates the strength of the proposed method, particularly when using the second-order Heun sampler during sampling. With a lower number of score function evaluations (NFE), our sampler achieves fast synthesis without compromising quality. We use 40 steps for sampling in all experiments, resulting in 119 NFEs with our settings. For comparison, we evaluate the Diffusion-TS using the same NFE. Given the trade-off of performance and sampling time in Diffusion-TS, our experimental results outperform previous methods. The marginal increase in time for TimeBridge is due to the time required to build the prior distribution.
@@
Table 6: Comparison of sampling efficiency and generation performance. The time column indicates the generation time of each sample in milliseconds (ms) with a batch size of 512.
Conclusion
In this paper, we highlight the effectiveness of selecting prior distributions for time series generation using diffusion bridges. To address the limitations in the standard Gaussian prior of the diffusion models, we propose a framework that allows for the selection of any prior suitable for various tasks. As a limitation, we leave experiments on other types of conditions, such as class labels or time-independent conditions, for future work. We believe that incorporating sampling guidance can address these types of conditions. As the proposed method can cover a wide range of scenarios in time series diffusion models, we believe this work significantly advances the investigation of prior distributions for developing a general time series generation framework.
@@ Appendix A Diffusion Bridge
Schrödinger Bridge.
Standard diffusion models suffer from limitations such as the requirement of a sufficiently long time for the prior distribution to approximate a standard Gaussian distribution and the computational expense associated with sample generation. To address these challenges, diffusion Schrödinger bridge models have been recently introduced (De Bortoli et al. 2021; Chen, Liu, and Theodorou 2021). The goal of the Schrödinger bridge problem is to find an entropic optimal transport between two probability distributions in terms of Kullback–Leibler divergence on path spaces within a finite horizon (Chen, Liu, and Theodorou 2021; Léonard 2012). While standard diffusion models and flow matching models are not guaranteed to provide optimal transport, Schrödinger bridge models aim to find paths that recover entropy-regularized versions of optimal transport (Shi et al. 2024). The formulation of the Schrödinger bridge problem is as follows:
$\underset{p\in\mathcal{P}_{[0,T]}}{min}D_{KL}(p\lVert p^{ref}),\quad s.t.\ p_{0}=p_{data},\ p_{T}=p_{prior},$ (15)
where $\mathcal{P}_{[0,T]}$ refers to the space of path measures on $[0,T]$, $p^{ref}$ denotes the reference path measure, and $p_{0},p_{T}$ represents the marginal distributions of $p$ at each time step $0,T$, respectively. The diffusion bridge algorithm is introduced by utilizing $p^{ref}$ from the forward SDE in Equation (1) (Vargas et al. 2021; Chen, Liu, and Theodorou 2021; De Bortoli et al. 2021). Setting the reference path as the forward SDE, the problem becomes equivalent to following forward-backward SDEs:
$\begin{split}d{\mathbf{x}}_{t}=[{\mathbf{f}}({\mathbf{x}}_{t},t)+g^{2}(t)\nabla\log\bm{\Psi}_{t}({\mathbf{x}}_{t})]dt+g(t)d{\mathbf{w}}_{t},\\ \quad{\mathbf{x}}_{0}\sim p_{data},\end{split}$ (16)
$\begin{split}d{\mathbf{x}}_{t}=[{\mathbf{f}}({\mathbf{x}}_{t},t)-g^{2}(t)\nabla\log\hat{\bm{\Psi}}_{t}({\mathbf{x}}_{t})]dt+g(t)d\bar{{\mathbf{w}}}_{t},\\ \quad{\mathbf{x}}_{T}\sim p_{prior},\end{split}$ (17)
@@
$\begin{split}\frac{\partial\bm{\Psi}}{\partial t}&=-\nabla_{x}\bm{\Psi}^{T}{\mathbf{f}}-\frac{1}{2}tr(g^{2}\nabla^{2}_{x}\bm{\Psi})\\ \frac{\partial\hat{\bm{\Psi}}}{\partial t}&=-\nabla_{x}\cdot(\hat{\bm{\Psi}}{\mathbf{f}})+\frac{1}{2}tr(g^{2}\nabla^{2}_{x}\hat{\bm{\Psi}})\\ &s.t.\quad\bm{\Psi}_{0}\hat{\bm{\Psi}}_{0}=p_{data},\bm{\Psi}_{T}\hat{\bm{\Psi}}_{T}=p_{prior},\\ &\qquad\quad p_{t}=\bm{\Psi}_{t}\hat{\bm{\Psi}}_{t}\ \text{for}\ t\in[0,T].\end{split}$ (18)
The Schrödinger bridge problem can be addressed using iterative methods (Kullback 1968; Fortet 1940). At the outset, several early studies employed iterative methods to simulate trajectories converging to the Schrödinger bridge problem (Vargas et al. 2021; Shi et al. 2024; De Bortoli et al. 2021). However, these methods rely on expensive iterative approximation techniques and have seen limited empirical application (Zhou et al. 2024).
Denoising Diffusion Bridge Models.
Alternatively, Chen, Liu, and Theodorou (2021) leveraged the theory of forward-backward SDEs to derive an exact log-likelihood expression for the Schrödinger bridge, which accurately generalizes the approach for score generative models. Drawing inspiration from the perspective that the framework of diffusion bridges, employing forward-backward SDEs, encompasses the paradigms of score matching diffusion models (Song, Meng, and Ermon 2020) and flow matching optimal transport paths (Lipman et al. 2022). Zhou et al. (2024) demonstrated that reintroducing several design choices from these domains becomes feasible. In particular, the reparameterization techniques outlined in Karras et al. (2022) are utilized.
@@
Table 7: VP and VE instantiations of diffusion bridge settings in DDBM.
For noise scheduling, common options include variance-preserving (VP) and variance-exploding (VE) diffusion (Song, Meng, and Ermon 2020). We express the reparameterization for the VP and VE bridge of Zhou et al. (2024) in Table 7. In our study, we choose to utilize the VP bridge, as our experimental findings indicate its superiority of VP over VE.
@@
$D_{\theta}({\mathbf{x}};\sigma)=c_{skip}(\sigma){\mathbf{x}}+c_{out}(\sigma)F_{\theta}(c_{in}(\sigma){\mathbf{x}};c_{noise}(\theta)),$ (19)
and trained $F_{\theta}$ instead of $D_{\theta}$ with the neural network. However, for time series data, estimating the diffusion noise poses greater challenges due to the presence of highly irregular noisy components (Shen and Kwok 2023; Yuan and Qiao 2024). Consistent with these studies, our experiments demonstrate that $D$-matching is more effective than $F$-matching. Detailed experimental results for comparison of VP with VE and $D$-matching with $F$-matching are provided in the ablation study in the Appendix D.
Appendix B Details of Datasets and Baselines
We first illustrate time series generation situations on various conditions in Figure 6.
@@ Appendix B Details of Datasets and Baselines
Figure 6: Illustration of time series generation situations on various conditions.
Data explanation.
@@
Table 8: Details of datasets.
Dataset Sample size # of features Link
@@
Baseline methods.
For the unconditional baseline methods, we compare with TimeVAE (Desai et al. 2021), TimeGAN (Yoon, Jarrett, and Van der Schaar 2019), Cot-GAN (Xu et al. 2020), Diffwave (Kong et al. 2021), DiffTime (Coletta et al. 2024), and Diffusion-TS (Yuan and Qiao 2024). For Table 1, we note the quality evaluation results in Diffusion-TS. We reimplement Diffusion-TS using their official GitHub in https://github.com/Y-debug-sys/Diffusion-TS, unfortunately, we cannot achieve the performance of Diffusion-TS in their original paper.
Furthermore, for the conditional generations, we reimplement SSSD (Alcaraz and Strodthoff 2023) using the official GitHub https://github.com/SSSD/sssd. For the trend-guided synthesis, we treat the trend as conditional observations and train the SSSD model to match the total time-series data. We add the conditional embedding layers in their original model architectures.
Appendix C Experimental settings
Model Architecture
We utilized Diffusion-TS (Yuan and Qiao 2024) as the based model, leveraging encoder-decoder transformer blocks. In this model, the encoder encodes the noisy sequence and the decoder consists of $K$ decoder blocks, where each block is composed of a transformer block, a feed-forward network, and interpretable decomposition layers to reflect the trends and seasonality of time series data. Each interpretable layers consist of trend and seasonal synthesis layers. $w_{(\cdot)}^{i,t}$ denotes the input of interpretable layers where $i\in 1,...,K$ denotes the index of the corresponding decoder block at diffusion step t.
We synthesize trend components using polynomial regressor as follows:
$V_{tr}^{t}=\Sigma_{i=1}^{K}(C\cdot\text{Linear}(w_{tr}^{i,t})+X_{tr}^{i,t}),$ (20)
where $X_{tr}^{i,t}$ is the mean value of the output of the $i^{th}$ decoder block, $\cdot$ denotes the tensor multiplication and $C=[1,c,...,c^{p}]\ \text{for}\ c=[0,1,2,...,\tau-1]^{T}/\tau$. We recover seasonal and error components using Fourier bases as follows:
$A_{i,t}^{(k)}=\lvert\mathcal{F}(w_{seas}^{i,t})_{k}\rvert,\Phi_{i,t}^{(k)}=\phi(\mathcal{F}(w_{seas}^{i,t})_{k}),$ (21)
$\kappa_{i,t}^{(1)},...,\kappa_{i,t}^{(K)}=\underset{k\in\{1,...,\lfloor\tau/2\rfloor+1\}}{\text{argTopK}}\{A_{i,t}^{(k)}\},$ (22)
$S_{i,t}=\Sigma_{k=1}^{K}A_{i,t}^{\kappa_{i,t}^{(k)}}[cos(2\pi f_{\kappa_{i,t}^{(k)}}\tau c+\Phi_{i,t}^{\kappa_{i,t}^{(k)}})+\\ cos(2\pi\bar{f}_{\kappa_{i,t}^{(k)}}\tau c+\bar{\Phi}_{i,t}^{\kappa_{i,t}^{(k)}})],$ (23)
where argTopK is to get the top $K$ amplitudes and K is a hyperparameter. $A_{i,t}^{(k)}$ and $\Phi_{i,t}^{(k)}$ are the phase, amplitude of the $k$-th frequency after the discrete Fourier transform $\mathcal{F}$, respectively. $f_{k}$ represents the Fourire frequency of the corresponding index $k$ and $\bar{(\cdot)}$ denotes ()($\cdot$)( ⋅ ) of the corresponding conjugates.
@@ Appendix C Experimental settings
We make details of the point-preserving sampler and its usage in imputation as proposed in Algorithm 1. We further explain the details of our sampler for the unconditional situation in Algorithm 2.
For implementing the diffusion sampler, we modify the official code of DDBM (Zhou et al. 2024) in https://github.com/alexzhou907/DDBM as the backbone of our sampler.
Input: Diffusion model $D_{\theta}({\mathbf{x}}_{t})$, score function ${\mathbf{s}}(D_{\theta},{\mathbf{x}}_{t_{i}},t_{i},{\mathbf{x}}_{T},{T})$ defined in Equation 9, diffusion steps $\{0=t_{0}<\cdots<t_{\Gamma}=T\}$; Condition ${\mathbf{y}}$, masking ${\mathbf{m}}$, step ratio $s$
Output: Denoised output ${\mathbf{x}}_{0}$
if Imputation  then ${\mathbf{x}}_{T}\leftarrow$ Linear spline interpolation using Equation (12) with masking ${\mathbf{m}}$
for $i=\Gamma,\ldots,1$ do
       Sample $\bm{{\bm{\varepsilon}}}\sim\mathcal{N}(\mathbf{0,I})$
       $\tilde{t}_{i}\leftarrow t_{i}+s(t_{i-1}-t_{i})$
       if $i\neq 1$ then
       end if
end for
Algorithm 1 Point-preserving Hybrid Sampler for TimeBridge
@@
As demonstrated in EDM (Karras et al. 2022), we can reparameterize $\alpha_{t}$ and $\sigma_{t}$ using $\beta_{min}$ and $\beta_{d}$ as follows:
$\alpha_{t}$ $=e^{-\frac{1}{4}\beta_{d}t^{2}-\frac{1}{2}\beta_{min}t}$ (24)
$\sigma_{t}$ $=e^{\frac{1}{2}\beta_{d}t^{2}+\beta_{min}t}-1,$ (25)
@@
Furthermore, we can adjust the weight scheduler of $w_{t}$ in Equation (13) with $\sigma_{data}$ as in EDM (Karras et al. 2022) and DDBM (Zhou et al. 2024) as follows:
$w_{t}=\frac{(\sigma_{t}^{2}+\sigma_{data}^{2})}{(\sigma_{t}\cdot\sigma_{data})^{2}}.$ (26)
Note that EDM (Karras et al. 2022) and DDBM (Zhou et al. 2024) uses this $\sigma_{data}$ in the reparmetriziation of $F$-matching in Equation (19) to calculate, $c_{skip},c_{out},c_{in},c_{noise}$, but we exclude them since we use direct $D$-matching.
@@
The summarized hyperparameters used for unconditional generation and conditional generation are demonstrated in Tables 9 and 10 for unconditional and conditional setting, respectively.
Table 9: Hyperparameters for unconditional generation.
@@
Base setup (Yuan and Qiao 2024) Batch size - 128 128 128 64 64 64
@@
Table 10: Hyperparameters for conditional generation.
@@
$\dagger$ We use 0.025 for the Energy dataset with Butterworth trend setting.

### v2

@@ 6.3. Synthesis with Trend Condition
We now consider trend-guided generation, where the data samples are in $\mathcal{D}=\{({\mathbf{x}}^{i},{\mathbf{y}}^{i})\}_{i=1}^{N}$. For the trend ${\mathbf{y}}$, we consider three types: linear, polynomial with degree of three, and Butterworth filter (Butterworth et al., 1930), a signal processing filter that captures trends (Xia et al., 2024). Using these trends as conditions, we reimplement the experiments in Table 2 with the ETTh and Energy datasets. To apply the conditions to baseline methods, we feed the trend into a conditional embedding and concatenate it with the original embedding. For TimeBridge, we set the prior ${\mathbf{x}}_{T}={\mathbf{y}}$ upon the conditional embedding.
@@
Experimental results are shown in Table 4. The rows with trend baseline indicate the results with the trends without any synthesis. After synthesis with trend conditions, we observe that the quality metrics diminished compared to the unconditional generation in Table 2, indicating that trend conditions help the model generate more realistic data samples. With the trend condition in ${\mathbf{x}}_{T}$, the proposed TimeBridge achieves the best performance in Context-FID in all settings and the best discriminative score in 5 out of 6 settings while maintaining correlation scores at similar levels to Diffusion-TS. This indicates that the proposed framework can be broadly adopted by different conditioning settings, without controlling any details other than the prior distribution. The sampling path of TimeBridge from trend to data is shown in Figure 4(a).
6.4. Imputation with Fixed Masks
@@
For imputation, we compare the proposed method with imputation diffusion models and the conditional Langevin sampler of Diffusion-TS (Yuan and Qiao, 2024). We use random masks for the Mujoco dataset and the geometric mask from (Zerveas et al., 2021) for the ETTh and Energy datasets. We set the total length of imputation to 100 for the Mujoco dataset, matching the data length, and split the ETTh and Energy datasets into 48 time steps. For the proposed TimeBridge, we preserve the observed values and set the prior ${\mathbf{x}}_{T}$ as the interpolated values of the condition ${\mathbf{c}}$ as in Equation (21).
@@
The imputation results are shown in Tables 5 and 6. Both tables demonstrate that TimeBridge achieves the lowest measures in imputation tasks. Since the Langevin sampler of Diffusion-TS does not require individual training for each missing ratio, we ensure a fair comparison by evaluating TimeBridge with the same training setting, TimeBridge-75%, which uses 75% missing values for training. TimeBridge-75% even outperforms the models trained with equal missing values, highlighting the generalization capability of TimeBridge in imputation tasks. The visualization of imputation results and sampling path can be found in Figures 5 and 4(b), respectively.
Figure 4. Illustration of sampling path from (Left) prior to (Right) data samples on the ETTh.
(a) ETTh Imputation
(b) Energy Imputation
@@
Figure 5. Visualization of imputation on (a) ETTh and (b) Energy dataset. The green dotted lines and circles denote true and observed values with their intervals, respectively.
6.5. Additional Experiments
@@
Table 7 evaluates the effectiveness of training design with three variants: (1) variance exploding (VE) scheduler, (2) noise matching instead of directly predicting $D_{\theta}$ as in Zhou et al. (2024), and (3) without Fourier-based loss during training. The results indicate that each component of our diffusion scheduler and loss design contributes to improved performance.
Extended ablation results are provided in Appendix C.6.
Further Experiments
Please refer to Appendix C for further experimental results and findings, including the impact of prior selection measured by Wasserstein distance, evaluations on extreme data scenarios, and results of predictive score (Yoon et al., 2019).
6.6. Computation efficiency
Table 8 illustrates the strength of the proposed method, particularly when using the second-order Heun sampler during sampling. With fewer score function evaluations (NFE), our sampler achieves fast synthesis without compromising quality. We use 40 steps for sampling in all experiments, resulting in 119 NFEs with our settings. For comparison, we evaluate the Diffusion-TS using the same NFE. Given the trade-off of performance and sampling time in Diffusion-TS, our experimental results outperform previous methods. The marginal increase in time for TimeBridge is due to the time required to build the prior distribution.
Table 7. Ablation study for training and sampling features.
@@ 6.6. Computation efficiency
Table 8. Comparison of sampling efficiency and generation performance. The time column indicates the generation time of each sample in milliseconds (ms) with a batch size of 512.
@@
7. Discussion and Conclusion
In this paper, we highlight the effectiveness of selecting prior distributions for time series generation using diffusion bridges. To address the limitations of the standard Gaussian prior in diffusion models, we propose a framework that allows the selection of any prior suitable for various tasks.
While the results are encouraging, several limitations remain. First, we rely on the standard architecture inherited from diffusion time-series synthesizers, which might not fully exploit the benefits of diffusion bridges. Future work should investigate transformer-based designs that incorporate prior information more effectively. Second, we leave the real-world applications with synthetic data beyond data synthesis itself for future study; realistic time-series synthesis could provide significant practical value.
As the proposed method can cover a wide range of scenarios in time series diffusion models, we believe this work significantly advances the investigation of prior distributions for developing a general time series generation framework. Ongoing advances in diffusion architectures will allow our approach to function as a plug-and-play adaptation that integrates with upcoming models.
@@ Appendix A Diffusion Bridge
A.1. Schrödinger Bridge
Standard diffusion models suffer from limitations such as the requirement of a sufficiently long time for the prior distribution to approximate a standard Gaussian distribution and the computational expense associated with sample generation. To address these challenges, diffusion Schrödinger bridge models have been recently introduced (De Bortoli et al., 2021; Chen et al., 2021). The goal of the Schrödinger bridge problem is to find an entropic optimal transport between two probability distributions in terms of Kullback–Leibler divergence on path spaces within a finite horizon (Chen et al., 2021; Léonard, 2012). While standard diffusion models and flow matching models are not guaranteed to provide optimal transport, Schrödinger bridge models aim to find paths that recover entropy-regularized versions of optimal transport (Shi et al., 2024). The formulation of the Schrödinger bridge problem is as follows:
(22) $\underset{p\in\mathcal{P}_{[0,T]}}{min}D_{KL}(p\lVert p^{ref}),\quad s.t.\ p_{0}=p_{data},\ p_{T}=p_{prior},$
where $\mathcal{P}_{[0,T]}$ refers to the space of path measures on $[0,T]$, $p^{ref}$ denotes the reference path measure, and $p_{0},p_{T}$ represents the marginal distributions of $p$ at each time step $0,T$, respectively. The diffusion bridge algorithm is introduced by utilizing $p^{ref}$ from the forward SDE in Equation (4) (Vargas et al., 2021; Chen et al., 2021; De Bortoli et al., 2021). Setting the reference path as the forward SDE, the problem becomes equivalent to following forward-backward SDEs:
(23) $\begin{split}d{\mathbf{x}}_{t}=[{\mathbf{f}}({\mathbf{x}}_{t},t)+g^{2}(t)\nabla\log\bm{\Psi}_{t}({\mathbf{x}}_{t})]dt+g(t)d{\mathbf{w}}_{t},\\ \quad{\mathbf{x}}_{0}\sim p_{data},\end{split}$
(24) $\begin{split}d{\mathbf{x}}_{t}=[{\mathbf{f}}({\mathbf{x}}_{t},t)-g^{2}(t)\nabla\log\hat{\bm{\Psi}}_{t}({\mathbf{x}}_{t})]dt+g(t)d\bar{{\mathbf{w}}}_{t},\\ \quad{\mathbf{x}}_{T}\sim p_{prior},\end{split}$
@@ A.1. Schrödinger Bridge
(25) $\begin{split}\frac{\partial\bm{\Psi}}{\partial t}&=-\nabla_{x}\bm{\Psi}^{T}{\mathbf{f}}-\frac{1}{2}tr(g^{2}\nabla^{2}_{x}\bm{\Psi})\\ \frac{\partial\hat{\bm{\Psi}}}{\partial t}&=-\nabla_{x}\cdot(\hat{\bm{\Psi}}{\mathbf{f}})+\frac{1}{2}tr(g^{2}\nabla^{2}_{x}\hat{\bm{\Psi}})\\ &s.t.\quad\bm{\Psi}_{0}\hat{\bm{\Psi}}_{0}=p_{data},\bm{\Psi}_{T}\hat{\bm{\Psi}}_{T}=p_{prior},\\ &\qquad\quad p_{t}=\bm{\Psi}_{t}\hat{\bm{\Psi}}_{t}\ \text{for}\ t\in[0,T].\end{split}$
The Schrödinger bridge problem can be addressed using iterative methods (Kullback, 1968; Fortet, 1940). At the outset, several early studies employed iterative methods to simulate trajectories converging to the Schrödinger bridge problem (Vargas et al., 2021; Shi et al., 2024; De Bortoli et al., 2021). A classical solver for Schrödinger bridges is Iterative Proportional Fitting (IPF). Beginning with arbitrary half-potentials $(\bm{\Psi}^{(0)}_{t},\hat{\bm{\Psi}}^{(0)}_{t})$, the algorithm alternates two normalization steps:
[…]
After each full iteration, the product $\bm{\Psi}^{(n+1)}_{t}\hat{\bm{\Psi}}^{(n+1)}_{t}$ matches both marginals, the sequence converges geometrically, requiring one forward–backward integration per iteration. However, these methods rely on expensive iterative approximation techniques and have seen limited empirical application (Zhou et al., 2024).
A.2. Denoising Diffusion Bridge Models
Alternatively, Chen et al. (2021) leveraged the theory of forward-backward SDEs to derive an exact log-likelihood expression for the Schrödinger bridge, which accurately generalizes the approach for score generative models. Drawing inspiration from the perspective that the framework of diffusion bridges, employing forward-backward SDEs, encompasses the paradigms of score matching diffusion models (Song et al., 2020) and flow matching optimal transport paths (Lipman et al., 2022). Zhou et al. (2024) demonstrated that reintroducing several design choices from these domains becomes feasible. In particular, the reparameterization techniques outlined in Karras et al. (2022) are utilized.
Table 9. VP and VE instantiations of diffusion bridge settings in DDBM.
@@ A.2. Denoising Diffusion Bridge Models
For noise scheduling, common options include variance-preserving (VP) and variance-exploding (VE) diffusion (Song et al., 2020). We express the reparameterization for the VP and VE bridge of Zhou et al. (2024) in Table 9. In our study, we choose to utilize the VP bridge, as our experimental findings indicate its superiority of VP over VE.
@@
(28) $D_{\theta}({\mathbf{x}};\sigma)=c_{skip}(\sigma){\mathbf{x}}+c_{out}(\sigma)F_{\theta}(c_{in}(\sigma){\mathbf{x}};c_{noise}(\theta)),$
and trained $F_{\theta}$ instead of $D_{\theta}$ with the neural network. However, for time series data, estimating the diffusion noise poses greater challenges due to the presence of highly irregular noisy components (Shen and Kwok, 2023; Yuan and Qiao, 2024). Consistent with these studies, our experiments demonstrate that $D$-matching is more effective than $F$-matching. Detailed experimental results for comparison of VP with VE and $D$-matching with $F$-matching are provided in the ablation study in the Appendix C.6.
Appendix B Experimental Settings
B.1. Situations
We illustrate time series generation situations on various conditions, including unconditional, linear trend, polynomial trend, fixed point, and missing value, in Figure 6.
@@ B.1. Situations
Figure 6. Illustration of time series generation situations under various conditions.
B.2. Baseline methods
For the unconditional baseline methods, we compare with TimeVAE (Desai et al., 2021), TimeGAN (Yoon et al., 2019), Cot-GAN (Xu et al., 2020), Diffwave (Kong et al., 2021), DiffTime (Coletta et al., 2024), and Diffusion-TS (Yuan and Qiao, 2024). For Table 2, we note the quality evaluation results in Diffusion-TS. We reimplement Diffusion-TS using their official GitHub in https://github.com/Y-debug-sys/Diffusion-TS, unfortunately, we cannot achieve the performance of Diffusion-TS in their original paper.
Furthermore, for the conditional generations, we reimplement SSSD (Alcaraz and Strodthoff, 2023) using the official GitHub https://github.com/SSSD/sssd. For the trend-guided synthesis, we treat the trend as conditional observations and train the SSSD model to match the total time series data. We add the conditional embedding layers in their original model architectures.
B.3. Evaluation Metrics
FID score is a widely used metric in image generation tasks, capturing the similarity between generated images and real ones (Heusel et al., 2017). To adapt it to the time series domain, Jeha et al. (2022) proposed Context-FID score, which replaces the Inception model used in the original FID with TS2Vec, a pre-trained time series representation learning model (Yue et al., 2022). As Context-FID score aligns with the predictive accuracy of the generative model in downstream tasks, a lower Context-FID score indicates higher synthetic data quality. For computation, first, we sample generated synthetic time series and real-time series separately. Next, both are encoded using a pre-trained TS2Vec model, and the FID score is calculated based on the resulting representations.
Correlational score is a metric used to evaluate the similarity in the covariance of time series features between original data and synthetic data (Liao et al., 2020). First, we estimate the covariance between the $i$-th and $j$-th features of the time series by:
(29) $cov_{i,j}=\frac{1}{T}\sum_{t=1}^{T}X_{i}^{t}X_{j}^{t}-\Big{(}\frac{1}{T}\sum_{t=1}^{T}X_{i}^{t}\Big{)}\Big{(}\frac{1}{T}\sum_{t=1}^{T}X_{j}^{t}\Big{)}.$
Then compute the correlation between the original data and synthetic data as follows:
(30) $\frac{1}{10}\sum_{i,j}^{d}\Big{|}\frac{cov_{i,j}^{o}}{\sqrt{cov_{i,i}^{o}cov_{j,j}^{o}}}-\frac{cov_{i,j}^{s}}{\sqrt{cov_{i,i}^{s}cov_{j,j}^{s}}}\Big{|},$
where $cov^{o}$ and $cov^{s}$ represent the covariance in the original data and synthetic data, respectively.
Discriminative score is a metric designed to evaluate the performance of a classifier in distinguishing between real and synthetic data. The score is calculated as the absolute difference between the classifier’s accuracy and 0.5.
B.4. Additional Experimental Details
Datasets
@@ B.4. Additional Experimental Details
Table 10. Details of datasets.
Dataset Sample size # of features Datatset link
@@
We further explain the details of our sampler for the unconditional situation in Algorithm 2. For implementing the diffusion sampler, we modify the official code of DDBM (Zhou et al., 2024) in https://github.com/alexzhou907/DDBM as the backbone of our sampler.
@@
As demonstrated in EDM (Karras et al., 2022), we can reparameterize $\alpha_{t}$ and $\sigma_{t}$ using $\beta_{min}$ and $\beta_{d}$ as follows:
(31) $\alpha_{t}$ $=e^{-\frac{1}{4}\beta_{d}t^{2}-\frac{1}{2}\beta_{min}t}$
(32) $\sigma_{t}$ $=e^{\frac{1}{2}\beta_{d}t^{2}+\beta_{min}t}-1,$
@@
Furthermore, we can adjust the weight scheduler of $w_{t}$ in Equation (18) with $\sigma_{data}$ as in EDM (Karras et al., 2022) and DDBM (Zhou et al., 2024) as follows:
(33) $w_{t}=\frac{(\sigma_{t}^{2}+\sigma_{data}^{2})}{(\sigma_{t}\cdot\sigma_{data})^{2}}.$
Note that EDM (Karras et al., 2022) and DDBM (Zhou et al., 2024) uses this $\sigma_{data}$ in the reparmetriziation of $F$-matching in Equation (28) to calculate, $c_{skip}$, $c_{out}$, $c_{in}$, $c_{noise}$, but we exclude them since we use direct $D$-matching.
@@
The summarized hyperparameters used for unconditional generation and conditional generation are demonstrated in Tables 11 and 12 for unconditional and conditional settings, respectively.
Table 11. Hyperparameters for unconditional generation.
@@
Base setup (Yuan and Qiao, 2024) Batch size - 128 128 128 64 64 64
@@
Table 12. Hyperparameters for conditional generation. $\dagger$ We use 0.025 for the Energy dataset with Butterworth trend setting.